# Non-linear conductances of Galton–Watson trees and application to the (near) critical random cluster model

### v1

@@ Abstract
When considering statistical mechanics models on trees, such that the Ising model, percolation, or more generally the random cluster model, some concave tree recursions naturally emerge. Some of these recursions can be compared with non-linear conductances, or $p$-conductances, between the root and the leaves of the tree. In this article, we estimate the $p$-conductances of $T_{n}$, a supercritical Galton–Watson tree of depth $n$, for any $p>1$ (for quenched realization of $T_{n}$). In particular, we find the sharp asymptotic behavior when $n$ goes to infinity, which depends on whether the offspring distribution admits a finite moment of order $q$, where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$. We then apply our results to the random cluster model on $T_{n}$ (with wired boundary condition) and provide sharp estimates on the probability that the root is connected to the leaves. As an example, for the Ising model on $T_{n}$ with plus boundary conditions on the leaves, we find that, at criticality, the quenched magnetization of the root decays like: (i) $n^{-1/2}$ times an explicit tree-dependent constant if the offspring distribution admits a finite moment of order $3$; (ii) $n^{-1/(\alpha-1)}$ if the offspring distribution has a heavy tail with exponent $\alpha\in(1,3)$.
@@ 1 Introduction and main results
We consider a super-critical branching process with reproduction law $\mu$, and we denote by $T$ the associated tree. We denote by $Z$ a generic random variable with law $\mu$ and we let $m:=\mathbf{E}[Z]$, that we assume to be finite. We assume for simplicity that $\mu(0)=0$ so the tree is infinite (and has no leaves) and that $\mu(1)<1$ so the tree is non-degenerate (and $m>1$).
For $n\in\mathbb{N}$, we let $T_{n}$ be the subtree of depth $n$, and we equip the tree $T_{n}$ with a set of resistances $R(e)$ on its edges: if the edge is $uv$ with $u$ the parent of $v$ (we will write $u\rightarrow v$), then we denote $R_{v}:=R(e)$. Our main objective is to estimate the $p$-conductance (or $p$-capacity) of $T_{n}$, equipped with resistances $(R_{v})_{v\in T_{n}}$. Such capacities arise naturally in the context of the random cluster model (including the percolation and Ising model) on a quenched Galton–Watson tree, we refer to Section 2 below for more details.
Let us mention that the usual effective resistance (or conductances) of random trees have already been considered in [1, 9], with a specific choice of (random) resistances $(R_{v})_{v\in T_{n}}$ which corresponds to a critical case. Our work can therefore be seen as a generalization of [1, 9] to the case of $p$-resistances, with a wider range of (non-random) resistances; we improve some of their results and present some applications, in particular to the (near) critical percolation or Ising models on a quenched tree.
More generally, we study (concave) recursions on a Galton–Watson tree: these recursions may arise in the context of statistical mechanics models on tree-like graphs (see [35] for an overview), or more generally in belief propagation or population dynamics algorithm (see e.g. [26] for an overview). These tree-recursion are often considered in a distributional sense rather than with a fixed tree, which corresponds to an annealed setting for random graphs (i.e. the randomness of the tree is part of the recursion); then, one of the main question is that of the convergence to the fixed point distribution, and we refer for instance to [20, 31] for recent (and general) results. Our interest is here slightly different, since we study iterations on a quenched (i.e. fixed) Galton–Watson tree, focusing on the case where the distributional fixed point of the iteration is degenerate, equal to 00: our main objective is then to estimate precisely the decay rate of the recursion towards zero.
@@ 1.1 Non-linear ($L^{p}$) resistive networks
Let us consider a graph $G=(V,E)$ equipped with a set of (non-negative) resistances $(R(e))_{e\in E}$ on its edges. A general theory of non-linear resistances and capacities is by now well-developed, and are usually defined through discrete nonlinear potential theory, see for instance [34] for an overview. Here, we focus on a specific non-linear case, so-called “$L^{p}$ resistive networks”. We will give the definitions directly in terms of the $L^{p}$-Thomson’s principle, since it is essentially the only tool we need for this article; we refer to [11] for a detailed review of $L^{p}$resistances and conductances, see in particular [11, Thm. 2.13] for the $L^{p}$-Thomson’s principle.
For $A,Z$ two disjoint subsets of $V$, we consider a flow $\theta$ between $A$ and $Z$, which is a function on ordered edges $\theta:E\to\mathbb{R}_{+}$ that verifies $\theta(-e)=-\theta(e)$ and Kirchoff’s node law: for $x\in V\setminus(A\cup Z)$, $\sum_{y\sim x}\theta(x,y)=0$. The strength of the flow is then
$\mathrm{Strength}(\theta)=\sum_{\begin{subarray}{c}x\in A,y\notin A\\ x\sim y\end{subarray}}\theta(x,y)$
@@
$\mathcal{E}_{p}(\theta)=\sum_{e\in E}R(e)^{\frac{1}{p-1}}\theta(e)^{\frac{p}{p-1}}\,.$
Then, the $p$-resistance and $p$-conductance (or $p$-capacity) between $A$ and $Z$ are defined, through Thomson’s principle as follows:
$\begin{split}\mathcal{R}_{p}(A\leftrightarrow Z)&:=\inf_{\begin{subarray}{c}\theta:A\to Z\\ \mathrm{Strength}(\theta)=1\end{subarray}}\mathcal{E}_{p}(\theta)^{p-1}\\ \text{ and }\quad\mathcal{C}_{p}(A\leftrightarrow Z)&:=\mathcal{R}_{p}(A\leftrightarrow Z)^{-1}\,.\end{split}$ (1.1)
@@
$\mathcal{R}_{p}(A\leftrightarrow Z)^{s}=\inf_{\begin{subarray}{c}\theta:A\to Z\\ \mathrm{Strength}(\theta)=1\end{subarray}}\hskip 2.84526pt\sum_{e\in E}R(e)^{s}\theta(e)^{q}\,,\qquad\text{ with }\ s:=q-1=\frac{1}{p-1}\,.$ (1.2)
Let us also stress that for $p=2$ ($q=2$), the $p$-resistance and $p$-conductance amount to the usual effective resistance and conductance.
@@
From the definition (1.1), we can easily deduce the Series and Parallel laws for $L^{p}$ electrical networks: we can formally write them as
@@
General non-linear networks can be defined by replacing the function $x\mapsto|x|^{p}$ by a strictly convex function $\varphi:x\mapsto\varphi(x)$ (in this case, Thomson’s principle becomes harder to state); we refer to [12, 34] and references therein. Most of our results would hold also for these capacities, but we have chosen to restrict to the case of $p$-capacities for simplicity, also because the applications we have in mind do not require more generality.
@@ 1.2 Concave recursions on trees
Let $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a concave function, and consider the iteration on the rooted tree $T_{n}$
$B_{n}(u)=\sum_{v\leftarrow u}\frac{R_{u}}{R_{v}}g\big{(}B_{n}(v)\big{)}\,,$ (1.3)
where we recall that $v\leftarrow u$ means that $v$ is a descendant of $u$ in $T_{n}$. These types of recursions appear in some statistical mechanic models on trees, in particular for the Random Cluster Model (with parameter $\mathtt{q}\in(0,2]$, $\mathtt{q}=1$ corresponding to the percolation model and $\mathtt{q}=2$ to the Ising model), see Section 2 for details. Let us now explain how they also appear when computing $L^{p}$ capacities of trees seen as resistive networks.
For $p>0$, and $u\in T_{n}\setminus\partial T_{n}$, let us consider the subtree $T_{n}(u)$ of all descendants of $u$ inside $T_{n}$, that we equip with resistances $(R_{u}R_{v})_{v\in T_{n}(u)}$; in other words, we simply multiply all original resistances inside $T_{n}(u)$ by $R_{u}$. We denote
$C_{n}^{(p)}(u):=\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$
the $p$-conductance from the root to the leaves of this subtree $T_{n}(u)$, as in (1.1). Then, the Series and Parallel laws from Remark 1.1 yield the following relation
$C_{n}^{(p)}(u)=\sum_{v\leftarrow u}\frac{R_{u}}{R_{v}}\frac{C_{n}^{(p)}(v)}{(1+C_{n}^{(p)}(v)^{s})^{1/s}}\,.$ (1.4)
This is exactly the iteration (1.3) with the function $g(x)=\frac{x}{(1+x^{s})^{1/s}}$ with $s:=q-1=\frac{1}{p-1}$, and boundary condition $C_{n}^{(p)}(u)=1$ if $u\in\partial T_{n}$.
@@
The recursion (1.4) appears in [33, Lem. 3.1], but with $s=p-1$ instead of $s=q-1$. This is just a convention, which simply means that all results in [33] are stated with the conjugate exponent; in practice $\mathrm{cap}_{p}$ in [33] should refer to the $\frac{p}{p-1}$-conductance.
We can actually compare (1.4) with (1.3) with a general concave $g$ if we make some assumptions on the function $g$. For instance, if we assume that $g$ is bounded and that there is some $q>1$ such that $x^{-q}(g(x)-x)$ remains bounded away from 00 and +$\infty$+ ∞ as $x\downarrow 0$, then we get that there are some constants $\kappa_{1}$, $\kappa_{2}$ such that for all $x>0$
$\frac{x}{(1+\kappa_{1}x^{s})^{1/s}}\leqslant g(x)\leqslant\frac{x}{(1+\kappa_{2}x^{s})^{1/s}}\,,$ (1.5)
with $s:=q-1$. This is for instance verified in the random cluster model of parameter $\mathtt{q}$ (with $q=2$ if $\mathtt{q}\in(0,2)$ and $q=3$ for $\mathtt{q}=2$), see Section 2 below.
One then gets the following proposition, similarly to Theorem 3.2 in [33].
@@
Let the function $g$ verify (1.5) for some $q>1$, and define $B_{n}(u)$ iteratively as in (1.3) with boundary condition $B_{n}(u)\in[c,+\infty]$ for $u\in\partial T$, for some constant $c>0$. Then, there are two constants $\kappa_{1}^{\prime}$ and $\kappa_{2}^{\prime}$ such that
$\kappa^{\prime}_{2}C_{n}^{(p)}(u)\leqslant B_{n}(u)\leqslant\kappa^{\prime}_{1}C_{n}^{(p)}(u)\,,$ (1.6)
where $p:=\frac{q}{q-1}$ is the conjugated exponent of $q$.
Let us also mention another interesting and easy lemma, which proves a monotonicity in $p$ of the conductances, see Lemma 6.9 in [3].
@@
For $1<p\leqslant p^{\prime}$, let $q$ (resp. $q^{\prime}$) be the conjugated exponent of $p$ (resp. $p^{\prime}$). Then, for any $u\in T_{n}$ we have $C_{n}^{(p)}(u)\geqslant C_{n}^{(p^{\prime})}(u)$. In particular, the $p$-conductance of a tree, $C_{n}^{(p)}(\rho)=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})$ is non-increasing in $p$, or equivalently non-decreasing in $q$.

### v2

@@ Abstract
In this article, we study concave recursions on trees, which appear widely in information theory through algorithms such as belief propagation, and in statistical mechanics through models on tree-like graphs, including the Ising model, percolation, and more generally, the random cluster model. These tree recursions can, in fact, be compared with non-linear conductances, or $p$-conductances, between the root and the leaves of the tree. In this article, we estimate the $p$-conductances of $T_{n}$, a supercritical Galton–Watson tree of depth $n$, for any $p>1$, for a quenched realization of $T_{n}$. In particular, we find the sharp asymptotic behavior when $n$ goes to infinity, which depends on whether the offspring distribution admits a finite moment of order $q$, where $q=\frac{p}{p-1}$ is the conjugate exponent of $p$. We then apply our results to the random cluster model on $T_{n}$ (with cluster weight parameter $\mathtt{q}\in(0,2]$ and wired boundary condition) providing sharp estimates on the probability that the root is connected to the leaves. As an example, for the Ising model on $T_{n}$ with plus boundary conditions on the leaves, we find that, at criticality, the quenched magnetization of the root decays like: (i) $n^{-1/2}$ times an explicit tree-dependent constant if the offspring distribution admits a finite third moment; (ii) $n^{-1/(\alpha-1)}$ if the offspring distribution has a heavy tail with exponent $\alpha\in(1,3)$.
@@ 1 Introduction and main results
We consider a super-critical branching process with reproduction law $\mu$. We denote by $Z$ a generic random variable with law $\mu$ and we denote $m:=\mathbf{E}[Z]$}, that we consider to be finite. We assume for simplicity that $\mu(0)=0$ so the tree is infinite (and has no leaves) and that $\mu(1)<1$ so the tree is non-degenerate (and $m>1$).
We denote by $T$ the infinite tree associated with the super-critical branching process, and for $n\in\mathbb{N}$, we let $T_{n}$ be the subtree of depth $n$. We equip the tree $T_{n}$ with a set of resistances $R(e)$ on its edges; if the edge is $uv$ with $u$ the parent of $v$ (we will write $u\rightarrow v$), then we denote $R_{v}:=R(e)$. Our main objective is to estimate the $p$-conductance (or $p$-capacity) of $T_{n}$, equipped with resistances $(R_{v})_{v\in T_{n}}$. Such quantities arise naturally in the context of the random cluster model with cluster weight $\mathtt{q}\in(0,2]$ (including percolation and the Ising model) on a quenched Galton–Watson tree; we refer to Section 2 below for more details (in particular, we also obtain a tree-recursion in the case $\mathtt{q}>2$, but it is not concave and cannot be related to a non-linear conductance).
Let us mention that the usual effective resistances (or conductances) of random trees have already been considered in [1, 10], with a specific choice of (random) resistances $(R_{v})_{v\in T_{n}}$ which corresponds to a critical case. Our work can therefore be seen as a generalization of [1, 10] to the case of $p$-resistances, with a wider range of (non-random) resistances; we improve some of their results and present some applications, in particular to the (near) critical percolation or Ising model on a quenched tree.
More generally, we study (concave) recursions on a Galton–Watson tree, which may arise in the context of statistical mechanics models on tree-like graphs (see [38] for an overview), or more broadly in belief propagation or population dynamics algorithms (see e.g. [29] for an overview). These tree-recursions are often considered in a distributional sense rather than with a fixed tree, corresponding to an annealed setting for random graphs (i.e. the randomness of the tree is part of the recursion); then, one of the main question is that of the convergence to the fixed point distribution, and we refer for instance to [22, 34] for recent (and general) results. Our interest here is slightly different, since we study iterations on a quenched (i.e. fixed) Galton–Watson tree, focusing on the case where the distributional fixed point of the iteration is degenerate, equal to 0: our main objective is then to estimate precisely the decay rate of the recursion towards zero. As an application of our results, we derive sharp estimates for the connection (or survival) probability in the sub-critical and (near) critical random cluster model with cluster weight $\mathtt{q}\in(0,2]$ on a quenched Galton–Watson tree.
@@ 1.1 Non-linear ($L^{p}$) resistive networks
Let us consider a graph $G=(V,E)$ equipped with a set of (non-negative) resistances $(R(e))_{e\in E}$ on its non-directed edges $e=\{x,y\}$ with $x,y\in V$. We will also need to consider directed edges, that are given by ordered couples $\vec{xy}:=(x,y)$ with $\{x,y\}\in E$; we write $\vec{e}=\vec{xy}$ and $\vec{e}=\vec{yx}$. A general theory of non-linear resistances and conductances is by now well-developed, and are usually defined through discrete nonlinear potential theory, see for instance [37] for an overview. Here, we focus on a specific non-linear case, so-called $L^{p}$ resistive networks”. We will give the definitions directly in terms of the $L^{p}$-Thomson’s principle, since it is the only tool we need for this article; we refer to [12] for a detailed review of $L^{p}$-resistances and conductances (or simply $p$-resistances and $p$-conductances), see in particular [12, Thm. 2.13] for the $L^{p}$-Thomson’s principle.
For $A,Z$ two disjoint subsets of $V$, we consider a flow $\theta$ between $A$ (the source) and $Z$ (the sink), which is a $\mathbb{R}$-valued function $\theta$ on directed edges that verifies $\theta(-\vec{e})=-\theta(\vec{e})$ and Kirchoff’s node law: for any $x\in V\setminus(A\cup Z)$, $\sum_{y:y\sim x}\theta(\vec{xy})=0$. The strength of the flow is then defined as
$\mathrm{Strength}(\theta)=\sum_{\begin{subarray}{c}x\in A,y\notin A\\ x\sim y\end{subarray}}\theta(\vec{xy})\,$,
@@
$\mathcal{E}_{p}(\theta)=\sum_{e\in E}R(e)^{\frac{1}{p-1}}|\theta(e)|^{\frac{p}{p-1}}\,$,
where $|\theta(e)|$ does not depend on the orientation of the edge since $\theta(-\vec{e})=-\theta(\vec{e})$. Then, the $p$-resistance and $p$-conductance between $A$ and $Z$ are defined, through Thomson’s principle, as follows:
$\begin{split}\mathcal{R}_{p}(A\leftrightarrow Z)&:=\inf_{\begin{subarray}{c}\theta:A\to Z\\ \mathrm{Strength}(\theta)=1\end{subarray}}\mathcal{E}_{p}(\theta)^{p-1}\\ \text{ and }\qquad\mathcal{C}_{p}(A\leftrightarrow Z)&:=\mathcal{R}_{p}(A\leftrightarrow Z)^{-1}\,.\end{split}$ (1.1)
@@
$\mathcal{R}_{p}(A\leftrightarrow Z)^{s}=\inf_{\begin{subarray}{c}\theta:A\to Z\\ \mathrm{Strength}(\theta)=1\end{subarray}}\hskip 2.84526pt\sum_{e\in E}R(e)^{s}|\theta(e)|^{q}\,,\qquad\text{ with }\ s:=q-1=\frac{1}{p-1}\,.$ (1.2)
Let us also stress that for $p=2$ (that is, $q=2$ and $s=1$), the $p$-resistance and $p$-conductance amount to the usual (linear) effective resistance and conductance.
@@
From the definition (1.1), we can easily deduce the Series and Parallel laws for $L^{p}$ resistive networks: recalling that $s=\frac{1}{p-1}$, we can formally write them as
@@
General non-linear networks can be defined by replacing the function $x\mapsto|x|^{p}$} by a strictly convex function $\varphi:x\mapsto\varphi(x)$ (in this case, Thomson’s principle becomes a bit harder to state); we refer to [13, 37] and references therein. Most of our results would also hold for such generalized conductances. We have chosen to restrict ourselves to the case of $p$-conductances for simplicity, and because the applications we have in mind do not require more generality.
@@ 1.2 Concave recursions on trees
The conductance recursion
It turns out that on trees, one can express the effective conductance between the root and the leaves $\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})$ through a simple recursion, as follows.
For $u\in T_{n}$, let us consider the subtree $T_{n}(u)$ of all descendants of $u$ inside $T_{n}$. We consider $\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$ the $p$-conductance of the subtree $T_{n}(u)$ as defined in (1.1), equal by convention to $\infty$ if $u\in\partial T_{n}$, and we denote
$C_{n}^{(p)}(u):=R_{u}\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))\,.$
Let us note that $C_{n}^{(p)}(u)$ corresponds to the $p$-conductance of the subtree $T_{n}(u)$ between its root $u$ and its leaves $\partial T_{n}(u)$, when equipped with resistances $(\frac{R_{v}}{R_{u}})_{v\in T_{n}(u)}$. Then, the Series and Parallel laws from Remark 1.1 easily yield the following relation
$C_{n}^{(p)}(u)=\sum_{v\leftarrow u}\frac{R_{u}}{R_{v}}\frac{C_{n}^{(p)}(v)}{(1+C_{n}^{(p)}(v)^{s})^{1/s}}\,,\quad\text{with }s:=\frac{1}{p-1}>0\,.$ (1.3)
@@
The recursion (1.3) appears in [36, Lem. 3.1], but with s=p1s=p-1 instead of $s=\frac{1}{p-1}$. This is just a convention, which simply means that all results in [36] are stated with the conjugate exponent; in practice the $p$-capacity in [36] (denoted by $\mathrm{cap}_{p}$) should refer to the $\frac{p}{p-1}$-conductance. Let us also mention that the term capacity is used in [36], but in our context it turns out to be equivalent to the effective conductance between the root and the leaves thanks to the Dirichlet and Thomson principles (see [12, Sec. 2, Thms. 2.10 and 2.14] where this is spelled out in the context of $L^{p}$ resistive networks). However, in other contexts, capacity might refer to the min-cut value, which typically differs from the conductance (except in trees). To avoid any ambiguity, we restrict ourselves to the term conductances.
General recursions and comparison argument
Consider the iteration on the rooted tree $T_{n}$ given by
$B_{n}(u)=\sum_{v\leftarrow u}f_{v}\big(B_{n}(v)\big)\,$, (1.4)
where $f_{v}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ are non-negative functions indexed by $v\in T_{n}$; recall also that $v\leftarrow u$ means that $v$ is a descendant of $u$ in $T_{n}$. Note that this is very similar to the iteration (1.3), if one could take $f_{v}(x)=\frac{R_{u}}{R_{v}}g_{s}(x)$ with the function $g_{s}(x):=\frac{x}{(1+x^{s})^{1/s}}$.
We will soon focus on the case where $f_{v}\equiv f$ does not depend on $v$, but let us keep the dependence on $v$ for now, to follow the footsteps of [36, §1.3]. We will further work with functions $(f_{v})_{v\in T_{n}}$ that are concave and bounded, and that satisfy the same type of expansion [^1] We use a standard notation: $\Theta$ is a function verifying $0<\liminf\frac{\Theta(\varepsilon)}{\varepsilon}\leqslant\limsup\frac{\Theta(\varepsilon)}{\varepsilon}<\infty$. $f_{v}(x)=K_{v}x-\Theta(x^{1+s})$ as $x\downarrow 0$, for some universal $s>0$ and some $K_{v}>0$ that may depend on $v$ — another formulation is through the comparison bounds (1.5) below.
Recursions like (1.8) appear in several statistical mechanics models on trees, in particular for the random cluster model with cluster weight $\mathtt{q}\in(0,2]$ (where $\mathtt{q}=1$ corresponds to the percolation model and $\mathtt{q}=2$ to the Ising model), we refer to Section 2 for details. They can actually be compared with $p$-conductances recursions on trees. Indeed, we can relate (1.4) to (1.3), provided that the functions $f_{v}$ can be suitably compared with $g_{s}$ and that the tree is equipped with appropriate resistances. The main comparison tool that we need for $f_{v}$ is the following, given by Theorem 3.2 in [36].
@@
Let $(f_{v})_{v\in T_{n}}$ be non-negative functions. Assume that there is some $s>0$ and some constants $\kappa_{1}$, $\kappa_{2}$ such that the following holds: for each $v\in T_{n}$ there is some $K_{v}>0$ such that for all $x>0$
$\frac{K_{v}x}{(1+\kappa_{1}x^{s})^{1/s}}\leqslant f_{v}(x)\leqslant\frac{K_{v}x}{(1+\kappa_{2}x^{s})^{1/s}}\,.$ (1.5)
Define $(B_{n}(u))_{u\in T_{n}}$ iteratively as in (1.4), with initial condition $B_{n}(u)=+\infty$ for $u\in\partial T_{n}$. Then, if we equip the tree $T_{n}$ with resistances $R_{v}:=\prod_{\rho\leqslant u\leqslant v}K_{u}^{-1}$, we have, for all $u\in T_{n}$
$\kappa_{1}^{-1/s}C_{n}^{(p)}(u)\leqslant B_{n}(u)\leqslant\kappa_{2}^{-2/s}C_{n}^{(p)}(u)\,,\qquad\text{with }p=\frac{1+s}{s}\,$, (1.6)
where we recall that $C_{n}^{(p)}(u):=R_{u}\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$.
Quick proof.
The proof simply proceed by iteration. For $\kappa>0$, let us define $F_{\kappa}(u)=\kappa^{-1/s}C_{n}^{(p)}(u)$, so that by (1.3) we have that for any $v\in T_{n}\setminus\partial T_{n}$,
$F_{\kappa}(v)=\sum_{v\leftarrow u}\frac{R_{u}}{R_{v}}\frac{F_{\kappa}(u)}{(1+\kappa F_{\kappa}(u)^{s})^{1/s}}\,.$
By definition we have $F_{\kappa_{1}}(u)=B_{n}(u)=F_{\kappa_{2}}(u)$ for all $u\in\partial T_{n}$. Then, we clearly get by iteration that $F_{\kappa_{1}}(u)\leqslant B_{n}(u)\leqslant F_{\kappa_{2}}(u)$, using (1.9) and the monotonicity of $x\mapsto\frac{x}{(1+\kappa x^{s})^{1/s}}$}. ∎
A similar proof establishes the monotonicity of the conductances in $p$, we state it in the following simple yet interesting lemma (taken from [3, Lemma 6.9]).
@@
Let $1<p\leqslant p^{\prime}$. Then, for any $u\in T_{n}$ we have $C_{n}^{(p)}(u)\geqslant C_{n}^{(p^{\prime})}(u)$. In particular, the $p$-conductance of a tree, $C_{n}^{(p)}(\rho)=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})$, is non-increasing in $p$, or equivalently non-decreasing in $s=\frac{1}{p-1}$.